If you stand 1 metre away from a certain speaker system turned to a particular song, the intensity that reaches your ears is 10−3 mW/m2, and the sound intensity level is 60 dB. If, instead you stand 10 metres away from the same system playing the same song, what do you expect will be the intensity of the sound?
1
Expert's answer
2020-02-10T09:30:17-0500
"I'=I\frac{r^2}{r'^2}"

"I'=10^{−3}\frac{1^2}{10^2}=10^{−5}\frac{mW}{m^2}=10^{−8}\frac{W}{m^2}"

Corresponding  sound intensity level is 40 dB.